OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (irishboy123):

\(y = -2x + 19 \\ y = x + 7\) so go for \(y = y\) ie: \( \ \ -2x + 19 = x + 7\) and see where you get

OpenStudy (anonymous):

y=11? and x=4?

OpenStudy (irishboy123):

\(\large −2x+19=x+7 \) \(\large −2x+19 (+ 2x)=x+7 (+ 2x)\) \(\large 19=3x+7\) \(\large 19-7=3x+7-7\) \(\large 12=3x\) \(\large 3x=12\) \(\large \frac{3x}{3}=\frac{12}{3}\) \(\large x = 4\) check: \(y = −2(4)+19 = 11\) and \(y =(4)+7\ = 11 \) \(\huge \checkmark \)

OpenStudy (anonymous):

awesome, thank you so much! i was a bit confused at first, you've helped greatly!